# From Heegaard diagrams to surgery

### v2

@@ 4. An example
The steps (S2) and (S3) in transforming our Heegaard diagram into a surgery are illustrated at the right-hand side of Figure 3. The step (S4) is illustrated by several pictures. At the left-hand side of Figure 4 we illustrate a parallel copy in $H$ of $\delta_{2}$, which is the curve denoted by 2 in our case. It is linked with the parallel copy of the curve denoted by 1 since it is pushed deeper into $H$. Then we make a parallel copy of $\delta_{3}$, which is the curve denoted by 3 in our case (see the right-hand side of Figure 4).
@@
Eventually, we obtain a surgery diagram illustrated in Figure 6. It is straightforward, by using Kirby’s calculus (see [6] and [4]), to transform this surgery into an unknot with framing 2, justifying that the manifold in question is $RP^{3}$.
Figure 6. The surgery diagram
Acknowledgements
We are grateful to Vladimir Grujić for pointing out to us the problem of surgery presentation of small covers and to Danica Kosanović for very useful comments concerning the previous version of this note. The first and the third author were supported by the Science Fund of the Republic of Serbia, Grant No. 7749891, Graphical Languages - GWORDS.

### v3

@@ 4. An example
The steps (S1) and (S2) in transforming our Heegaard diagram into a surgery are illustrated in Figure 7. In the step (S3) each curve at the right-hand side of Figure 7, except the blue one whose framing is 3, obtains the framing equal to 1. However, we will keep the numbers of curves until the last figure, where the framing is present. This is done in order to help the reader to follow the example. The step (S4) is illustrated by several pictures. At the left-hand side of Figure 8 we illustrate a parallel copy in $H$ of $\delta_{2}$, which is the curve denoted by 2 in our case. It is linked with the parallel copy of the curve denoted by 1 since it is pushed deeper into $H$. Then we make a parallel copy of $\delta_{3}$, which is the curve denoted by 3 in our case (see the right-hand side of Figure 8).
@@
Eventually, we obtain a surgery diagram (with framing) illustrated in Figure 10. It is straightforward, by using Kirby’s calculus (see [7] and [5]), to transform this surgery into an unknot with framing 2, justifying that the manifold in question is $RP^{3}$.
Figure 10. A surgery diagram for $RP^{3}$
Acknowledgements. We are grateful to the anonymous referee for helping us to improve the paper. We are also grateful to Vladimir Grujić for pointing out to us the problem of surgery presentation of small covers and to Danica Kosanović for very useful comments concerning the previous version of this note.
Declarations
Funding J.N. and Z.P. were supported by the Science Fund of the Republic of Serbia, Grant No. 7749891, Graphical Languages - GWORDS.
Competing interests The authors have no relevant financial or non-financial interests to disclose.